# Fast post-process Bayesian inference with Variational Sparse Bayesian Quadrature

### v3

@@ 1 Introduction
Bayesian inference is a well-founded approach to uncertainty quantification and model selection, widely adopted in data science and machine learning (Robert et al., 2007; Gelman et al., 2013; Ghahramani, 2015). Key quantities in Bayesian inference are the posterior distribution of model parameters, useful for parameter estimation and uncertainty quantification, and the marginal likelihood or model evidence, useful for model selection. In practice, Bayesian inference is particularly challenging when dealing with models with ‘black-box’ features common in science and engineering, such as lack of gradients or mildly-to-very expensive and possibly noisy evaluations of the likelihood, e.g., arising from simulator-based estimation (Wood, 2010; van Opheusden et al., 2020).
@@
Related work.
There are several works related to post-processing existing density evaluations to construct an approximate posterior. Zhang et al. (2022) find the best multivariate normal approximation along the optimization paths generated by a quasi-Newton optimization algorithm, L-BFGS (Liu and Nocedal, 1989), in terms of Kullback-Leibler (KL) divergence to the true posterior. Unlike our method, the L-BFGS algorithm requires the gradients of the log-likelihood or log-density function to build approximate estimates of the Hessian along the optimization trajectory and is brittle to noise in the target function. Moreover, the post-process estimation of the KL divergence needs additional evaluations of the log-density function. Yao et al. (2022) propose to reuse parallel – and possibly incomplete – runs of different inference algorithms such as Markov Chain Monte Carlo (mcmc; Robert and Casella, 2004, see e.g.) or variational inference (Blei et al., 2017) by combining them in a weighted average via ‘Bayesian stacking’. Similar in spirit to our approach, the computation of the stacking weights only requires a post-processing step. However, Bayesian stacking requires the samples to be approximately drawn from the posterior and does not make use of all available target evaluations. Finally, a related technique that merits mention is offline black-box optimization (Krishnamoorthy et al., 2023; Trabucco et al., 2022). In offline black-box optimization, the objective is to identify a parameter input that maximizes a black-box function utilizing pre-existing offline function evaluations. Much like our approach, offline black-box optimization capitalizes on leveraging existing evaluations. However, a key distinction lies in our primary objective, which is to construct an approximation of the unknown posterior density, a task inherently more challenging than finding a point estimate.
Outline.
@@ 2 Background
In this section, we present the core concepts and techniques used in the paper. We recall that our objective is to compute a tractable approximation of the posterior density, given the evaluations (observations) of the unnormalized log-density function. We provide an overview of key techniques including the variational inference method for posterior approximation, the Gaussian process as a regression surrogate, Bayesian quadrature, Variational Bayesian Monte Carlo (Acerbi, 2018), and sparse Gaussian processes.
Notation.
Throughout the paper, we denote with $f_{0}(\mathbf{x})\equiv\log p(\mathcal{D}|\mathbf{x})p(\mathbf{x})$ the target (unnormalized) log posterior density or log joint, where $p(\mathcal{D}|\mathbf{x})$ is the likelihood of the model of interest under the data $\mathcal{D}$, $p(\mathbf{x})$ is the prior, and $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{D}$ is a vector of model parameters. The dimension of the parameter space is denoted as $D\in\mathbb{N}$. We indicate with $(\mathbf{x}_{n},y_{n})$ pairs of observed locations and values of the log-density, i.e., $y_{n}=f_{0}(\mathbf{x}_{n})$. For noisy evaluations of the target arising from simulator-based estimates, $\sigma_{\text{obs}}^{2}(\mathbf{x}_{n})$ denotes the variance of the observation, assumed to be known (in practice, estimated; for example see Acerbi, 2020). $\left(\mathbf{X},\mathbf{y},\mathbf{s}\right)\equiv\{\mathbf{x}_{n},y_{n},\sigma_{\text{obs}}(\mathbf{x}_{n})\}_{n=1}^{N}$ denote the set of observations. We denote with $\mathbf{S}$ the diagonal matrix of observation noise variances, with $\mathbf{S}\equiv\text{diag}\left[\sigma_{\text{obs}}^{2}(\mathbf{x}_{1}),\ldots,\sigma_{\text{obs}}^{2}(\mathbf{x}_{N})\right]$. [^2] Later, this will be the total observation noise, which includes noise shaping introduced in Section 3.3.
@@ 2.1 Variational inference
Variational inference is a popular approach to approximate the intractable posterior density $p(\mathbf{x}|\mathcal{D})$ via a simpler distribution $q(\mathbf{x})\equiv q_{\bm{\phi}}(\mathbf{x})$ that belongs to a parametric family indexed by $\bm{\phi}$ (Jordan et al., 1999; Bishop, 2006; Blei et al., 2017). The goal of variational inference is to find $\bm{\phi}$ for which the variational posterior $q_{\bm{\phi}}$ best approximates the true posterior, as quantified by the reverse Kullback-Leibler (KL) divergence,
@@
where $\mathbb{E}_{\bm{\phi}}\equiv\mathbb{E}_{q_{\bm{\phi}}}$. Crucially, $D_{\text{KL}}(q||p)\geq 0$ and $D_{\text{KL}}(q||p)=0$ if and only if $q=p$. Minimizing the KL divergence casts Bayesian inference as an optimization problem. This optimization consists of finding the variational parameters, $\bm{\phi}$, that maximize the objective:
$\begin{split}\text{{elbo}}\left[q_{\bm{\phi}}\right]&=\mathbb{E}_{\bm{\phi}}\left[\log\frac{p(\mathcal{D}|\mathbf{x})p(\mathbf{x})}{q_{\bm{\phi}}(\mathbf{x})}\right]\\ &=\mathbb{E}_{\bm{\phi}}\left[f_{0}(\mathbf{x})\right]+\mathcal{H}[q_{\bm{\phi}}(\mathbf{x})],\end{split}$ (2)
@@ 2.2 Gaussian processes
When a computational model of interest has an expensive black-box likelihood, a proven approach for efficient Bayesian inference consists of building a (cheaper) surrogate model to emulate either the log-likelihood $\log p(\mathcal{D}|\mathbf{x})$ or directly the log joint density function $f_{0}(\mathbf{x})$; see Figure 1. There is a long tradition of using Gaussian processes (GPs) as surrogate models for Bayesian inference (Rasmussen, 2003; Gunter et al., 2014; Gutmann and Corander, 2016; Nemeth and Sherlock, 2018; Wang and Li, 2018; Acerbi, 2018; Järvenpää et al., 2021; De Souza et al., 2022; Gammal et al., 2022).
@@
Tractable surrogate density approximation.
A key observation is that a GP surrogate of the log density (Figure 1a) does not immediately yield a usable approximate posterior. The normalization constant is unknown and we cannot directly sample from the density associated with the GP surrogate. Another layer of approximation is needed to go from the log density surrogate to posterior approximation (Figure 1). Several approximation techniques have been developed to tackle this issue, many of them involving Bayesian quadrature (Ghahramani and Rasmussen, 2002b; Osborne et al., 2012; Gunter et al., 2014; Acerbi, 2018; Adachi et al., 2022).
@@ 2.3 Bayesian quadrature
Many key computations in Bayesian inference require the estimation of intractable integrals, as for example the elbo seen in Eq. 2. Bayesian quadrature (O’Hagan, 1991; Ghahramani and Rasmussen, 2002a) is a technique to obtain Bayesian estimates of intractable integrals of the form
@@
where $f$ is a function of interest and $\pi$ a known probability distribution. Here we consider the domain of integration $\mathcal{X}=\mathbb{R}^{D}$. When a GP prior is specified for $f$, since integration is a linear operator, the integral $\mathcal{J}$ is also a Gaussian random variable whose posterior mean and variance are (Ghahramani and Rasmussen, 2002a)
@@ 2.4 Variational Bayesian Monte Carlo
$\mathbb{E}_{f}\left[\text{{elbo}}(\bm{\phi})\right]=\mathbb{E}_{f}\left[\mathbb{E}_{\bm{\phi}}\left[f\right]\right]+\mathcal{H}[q_{\bm{\phi}}],$ (8)
@@ 2.5 Sparse Variational Gaussian Processes
Sparse Gaussian process regression.
@@
For the case of sparse GP regression with Gaussian likelihood, Titsias (2009) proved that the optimal variational distribution that maximizes Eq. 12 for fixed GP hyperparameters and inducing point locations $\mathbf{Z}$ is given by $\tilde{p}(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}})$, defined below. Furthermore, equality between the sparse and full GP posterior is obtained for $\mathbf{Z}=\mathbf{X}$. The optimal variational parameters $\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}}$ of the sparse GP are:
@@
where $\bm{\Sigma}\equiv(\mathbf{K}_{\mathbf{Z},\mathbf{X}}\mathbf{\mathbf{S}}^{-1}\mathbf{K}_{\mathbf{X},\mathbf{Z}}+\mathbf{K}_{\mathbf{Z},\mathbf{Z}})^{-1}$.
Detailed derivations and numerical implementation details are provided in Supplementary Material A.1-A.3.
@@ 3.1 Overview of the algorithm
The output of the procedure is an approximate posterior $q_{\phi}$, the estimated surrogate elbo (see Section 2.4), and the surrogate elbo’s standard deviation $\textsc{elbo}_{\text{sd}}$. We briefly describe each step below and discuss additional details in the following sections.
Trimming of the provided evaluations.
We remove from the provided log-density evaluations all points with extremely low log-density values relative to the maximum observed value. Such points are at best weakly informative (indication of near-zero density value) for approximating the posterior and at worst induce instabilities in the GP surrogate (Acerbi et al., 2018; Järvenpää et al., 2021; Gammal et al., 2022). Note that we keep points with low density values, as those are useful for anchoring the GP surrogate (De Souza et al., 2022) – we just remove the extremely low ones, as explained below. The remaining points after removal are typically still of the order of many thousands, too many to be handled efficiently by exact GPs. We call this preprocessing step trimming.
@@
$\mu_{\bm{\psi}}(\tilde{\mathbf{x}})$ $=\kappa(\tilde{\mathbf{x}},\mathbf{Z})\bm{\Sigma}\mathbf{K}_{\mathbf{Z},\mathbf{X}}\mathbf{S}^{-1}\mathbf{y},$ (18)
@@
where $\bm{\Sigma}$ is defined in Eq. 14.
@@
Sparse Bayesian quadrature.
@@
$\begin{split}\mathbb{E}\left[\mathcal{I}_{k}\right]=&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{k},\bm{\Sigma}_{k})f(\tilde{\mathbf{x}}\mid\textbf{u})\psi(\textbf{u})\mathrm{d}\tilde{\mathbf{x}}\mathrm{d}\textbf{u}\\ =&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{k},\bm{\Sigma}_{k})\mu_{\bm{\psi}}(\tilde{\mathbf{x}})\mathrm{d}\tilde{\mathbf{x}},\\ \end{split}$ (22)
@@ 3.2 Inducing points and hyperparameter selection
Inducing point location.
@@
For known GP hyperparameters, Burt et al. (2020) show that sampling from a $M$-determinantal point process ($M$-DPP) to select inducing points $\mathbf{Z}\subset\mathbf{X}$ will make the trace error term of the GP-elbo (Eq. 24) close to its optimal value. Since the DPP approach is intractable, we follow Burt et al. (2020) and Maddox et al. (2021) who recommend instead to sequentially select inducing points that greedily maximize the diagonal of the error term,
@@
GP hyperparameters.
The procedure for the selection of inducing points mentioned above requires known GP hyperparameters $\bm{\xi}$. In practice, it is often enough to start with a reasonable estimate for $\bm{\xi}$ and then iterate the process multiple times (Burt et al., 2020). To obtain an initial estimate for the GP hyperparameters, we fit an exact GP on a small subset of the data, and use the exact GP hyperparameters for initial inducing point selection. The subset is chosen via stratified $K$-means, where we ensure that the chosen subset is representative of the full set in terms of both location and log-density values. After the above initialization, we iterate sparse GP hyperparameter optimization and inducing points selection using the current sparse GP hyperparameters, until no improvement on the GP-elbo is found. Similarly to other block-optimization procedures, this process is not guaranteed to find the global optimum of the GP-elbo, but it often works well in practice (Burt et al., 2020).

### v4

@@ 1 Introduction
Bayesian inference is a well-founded approach to uncertainty quantification and model selection, widely adopted in data science and machine learning (Robert et al., 2007; Gelman et al., 2013; Ghahramani, 2015). Key quantities in Bayesian inference are the posterior distribution of model parameters, useful for parameter estimation and uncertainty quantification, and the marginal likelihood or model evidence, useful for model selection. In practice, Bayesian inference is particularly challenging when dealing with models with ‘black-box’ features common in science and engineering, such as lack of gradients or mildly-to-very expensive and possibly noisy evaluations of the likelihood, e.g., arising from simulation-based estimation (Diggle and Gratton, 1984; Wood, 2010; van Opheusden et al., 2020; Price et al., 2018a).
@@
1.1 Related work
There are several works related to post-processing existing density evaluations to construct an approximate posterior. Zhang et al. (2022) find the best multivariate normal approximation along the optimization paths generated by a quasi-Newton optimization algorithm, L-BFGS (Liu and Nocedal, 1989), in terms of Kullback-Leibler (KL) divergence to the true posterior. Unlike our method, the L-BFGS algorithm requires the gradients of the log-likelihood or log-density function to build approximate estimates of the Hessian along the optimization trajectory and is brittle to noise in the target function. More closely aligned with our approach, Bliznyuk et al. (2008) locates the high posterior density region through derivative-free MAP optimization and augments the evaluation set with additional design points to build a surrogate of the log density using radial basis functions. Both methods above require additional evaluations of the log-density function. Yao et al. (2022) propose to reuse parallel – and possibly incomplete – runs of different inference algorithms such as Markov Chain Monte Carlo (mcmc; Robert and Casella, 2004, see e.g.) or variational inference (Blei et al., 2017) by combining them in a weighted average via ‘Bayesian stacking’. Similar in spirit to our approach, the computation of the stacking weights only requires a post-processing step. However, Bayesian stacking requires the samples to be approximately drawn from the posterior and does not make use of all available target evaluations.
Our work also connects to simulation-based inference (SBI), a broad framework for estimating posteriors when likelihoods are intractable or computationally expensive (Diggle and Gratton, 1984; Cranmer et al., 2020). SBI methods include classical approaches such as rejection sampling (Sisson et al., 2018), parametric and nonparametric likelihood approximations (Diggle and Gratton, 1984; Price et al., 2018b; Gutmann and Corander, 2016b), and modern neural density estimation techniques (Lueckmann et al., 2021; Radev et al., 2020). Our approach is distinct in that it post-processes a fixed set of (noisy) log-density evaluations to estimate a single posterior, without requiring additional simulations. Related to the post-processing idea, Yao et al. (2024) proposes to improve posterior approximations via stacking multiple SBI results.
Finally, a related technique that merits mention is offline black-box optimization (Krishnamoorthy et al., 2023; Trabucco et al., 2022). In offline black-box optimization, the objective is to identify a parameter input that maximizes a black-box function utilizing pre-existing offline function evaluations. Much like our approach, offline black-box optimization capitalizes on leveraging existing evaluations. However, a key distinction lies in our primary objective, which is to construct an approximation of the unknown posterior density, a task inherently more challenging than finding a point estimate.
1.2 Outline
@@ 2 Background
In this section, we present the core concepts and techniques used in the paper. We recall that our objective is to compute a tractable approximation of the posterior density, given the evaluations (observations) of the unnormalized log-density function. We provide an overview of key techniques, including the variational inference method for posterior approximation, the Gaussian process as a regression surrogate, Bayesian quadrature, Variational Bayesian Monte Carlo (Acerbi, 2018), and sparse Gaussian processes.
2.1 Notation
Throughout the paper, we denote with $f_{0}(\mathbf{x})\equiv\log p(\mathcal{D}|\mathbf{x})p(\mathbf{x})$ the target (unnormalized) log posterior density or log joint, where $p(\mathcal{D}|\mathbf{x})$ is the likelihood of the model of interest under the data $\mathcal{D}$, $p(\mathbf{x})$ is the prior, and $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{D}$ is a vector of model parameters. The dimension of the parameter space is denoted as $D\in\mathbb{N}$. We indicate with $(\mathbf{x}_{n},y_{n})$ pairs of observed locations and values of the log-density, i.e., $y_{n}=f_{0}(\mathbf{x}_{n})$. For noisy evaluations of the target arising from simulation-based estimates, $\sigma_{\text{obs}}^{2}(\mathbf{x}_{n})$ denotes the variance of the observation, assumed to be known—in practice, it is typically estimated (see, for example, Acerbi, 2020; Järvenpää et al., 2021). $\left(\mathbf{X},\mathbf{y},\mathbf{s}\right)\equiv\{\mathbf{x}_{n},y_{n},\sigma_{\text{obs}}(\mathbf{x}_{n})\}_{n=1}^{N}$ denote the set of observations. We denote with $\mathbf{S}$ the diagonal matrix of observation noise variances, with $\mathbf{S}\equiv\text{diag}\left[\sigma_{\text{obs}}^{2}(\mathbf{x}_{1}),\ldots,\sigma_{\text{obs}}^{2}(\mathbf{x}_{N})\right]$. [^2] Later, this will be the total observation noise, which includes noise shaping introduced in Section 3.3.
@@ 2.2 Variational inference
Variational inference is a popular approach to approximate the intractable posterior density $p(\mathbf{x}|\mathcal{D})$ via a simpler distribution $q(\mathbf{x})\equiv q_{\bm{\phi}}(\mathbf{x})$ that belongs to a parametric family indexed by $\bm{\phi}$ (Jordan et al., 1999; Bishop, 2006; Blei et al., 2017; Kingma and Welling, 2013). The goal of variational inference is to find $\bm{\phi}$ for which the variational posterior $q_{\bm{\phi}}$ best approximates the true posterior, as quantified by the reverse Kullback-Leibler (KL) divergence,
@@
where $\mathbb{E}_{\bm{\phi}}\equiv\mathbb{E}_{\mathbf{x}\sim q_{\bm{\phi}}(\mathbf{x})}$. Crucially, $D_{\text{KL}}(q||p)\geq 0$ and $D_{\text{KL}}(q||p)=0$ if and only if $q=p$. Minimizing the KL divergence casts Bayesian inference as an optimization problem. This optimization consists of finding the variational parameters, $\bm{\phi}$, that maximize the objective:
$\begin{split}\text{{elbo}}\left(q_{\bm{\phi}}\right)&=\mathbb{E}_{\bm{\phi}}\left[\log\frac{p(\mathcal{D}|\mathbf{x})p(\mathbf{x})}{q_{\bm{\phi}}(\mathbf{x})}\right]\\ &=\mathbb{E}_{\bm{\phi}}\left[f_{0}(\mathbf{x})\right]+\mathcal{H}[q_{\bm{\phi}}(\mathbf{x})],\end{split}$ (2)
@@ 2.3 Gaussian processes
When a computational model of interest has an expensive black-box likelihood, a proven approach for efficient Bayesian inference consists of building a (cheaper) surrogate model to emulate either the log-likelihood $\log p(\mathcal{D}|\mathbf{x})$ or directly the log joint density function $f_{0}(\mathbf{x})$; see Figure 1. There is a long tradition of using Gaussian processes (GPs) as surrogate models for Bayesian inference (Rasmussen, 2003; Gunter et al., 2014; Gutmann and Corander, 2016a; Nemeth and Sherlock, 2018; Wang and Li, 2018; Acerbi, 2018; Järvenpää et al., 2021; De Souza et al., 2022; El Gammal et al., 2023).
@@
Tractable surrogate density approximation
A key observation is that a stochastic GP surrogate of the log density (Figure 1a) does not immediately yield a usable approximate posterior. The normalization constant is unknown and we cannot directly sample from the density associated with the GP surrogate. Another layer of approximation is needed to go from the log density surrogate to posterior approximation (Figure 1). One straightforward approach is to use the posterior mean function of the GP as a deterministic surrogate for the log density and apply MCMC methods to sample from the resulting approximate posterior (Järvenpää et al., 2021; El Gammal et al., 2023). Alternatively, several techniques leverage Bayesian quadrature (O’Hagan, 1991) to solve the integrals involving the stochastic GP surrogate (Rasmussen and Ghahramani, 2002; Osborne et al., 2012; Gunter et al., 2014; Acerbi, 2018; Adachi et al., 2022), as described next.
@@ 2.4 Bayesian quadrature
Many key computations in Bayesian inference require the estimation of intractable integrals, for example the elbo seen in Eq. 2. Bayesian quadrature (O’Hagan, 1991; Rasmussen and Ghahramani, 2002) is a technique to obtain Bayesian estimates of intractable integrals of the form
@@
where $f$ is a function of interest and $\pi$ a known probability distribution. Here we consider the domain of integration $\mathcal{X}=\mathbb{R}^{D}$. When a GP prior is specified for $f$, since integration is a linear operator, the integral $\mathcal{J}$ is also a Gaussian random variable whose posterior mean and variance are (Rasmussen and Ghahramani, 2002)
@@ 2.5 Variational Bayesian Monte Carlo
$\overline{\textsc{elbo}}=\mathbb{E}_{f}\left[\text{{elbo}}(q_{\bm{\phi}})\right]=\mathbb{E}_{f}\left[\mathbb{E}_{\bm{\phi}}\left[f\right]\right]+\mathcal{H}[q_{\bm{\phi}}],$ (8)
@@ 2.6 Sparse Variational Gaussian Processes
Sparse Gaussian process regression
@@
For the case of sparse GP regression with Gaussian likelihood, Titsias (2009) proved that the optimal variational distribution that maximizes Eq. 12 for fixed GP hyperparameters and inducing point locations $\mathbf{Z}$ is given by $\tilde{p}(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}})$, with
@@
where $\bm{\Sigma}\equiv(\mathbf{K}_{\mathbf{Z},\mathbf{X}}\mathbf{\mathbf{S}}^{-1}\mathbf{K}_{\mathbf{X},\mathbf{Z}}+\mathbf{K}_{\mathbf{Z},\mathbf{Z}})^{-1}$. We denote by $\bm{\psi}=(\mathbf{m}_{\mathbf{u}},\mathbf{R}_{\mathbf{u}\mathbf{u}})$ the optimal variational parameters and $p(\mathbf{f},\mathbf{u}\mid\bm{\psi})=q(\mathbf{f},\mathbf{u})$ the joint variational posterior of the sparse GP under this optimal setting. Note that equality between the sparse and full GP posterior is obtained for $\mathbf{Z}=\mathbf{X}$.
Detailed derivations for the non-zero mean case and numerical implementation details are provided in Supplementary Material A.1-A.3.
@@ 3.1 Overview of the algorithm
The output of the procedure is an approximate posterior $q_{\phi}$, along with the estimated surrogate elbo mean $\overline{\textsc{elbo}}$ (see Section 2.5) and its standard deviation $\textsc{elbo}_{\text{sd}}$. We briefly describe each step below and discuss additional details in the following sections. A complete algorithmic description is available in Supplementary Material C.1.
Trimming of the provided evaluations
We remove from the provided log-density evaluations all points with extremely low log-density values relative to the maximum observed value. Such points are at best weakly informative (indication of near-zero density value) for approximating the posterior and at worst induce instabilities in the GP surrogate (Acerbi et al., 2018; Järvenpää et al., 2021; El Gammal et al., 2023). Note that we keep points with low density values, as those are useful for anchoring the GP surrogate (De Souza et al., 2022) – we just remove the extremely low ones, as explained below. The remaining points after removal are typically still of the order of many thousands, too many to be handled efficiently by exact GPs. We call this preprocessing step trimming.
@@
$\mu_{\bm{\psi}}(\tilde{\mathbf{x}})=$ $\kappa(\tilde{\mathbf{x}},\mathbf{Z})\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-1}(\mathbf{m}_{\mathbf{u}}-m(\mathbf{Z}))+m(\tilde{\mathbf{x}}),$ (18)
@@
where $\mathbf{m}_{\mathbf{u}}$ and $\bm{\Sigma}$ are defined in Eqs. 13 and 14.
@@
Sparse Bayesian quadrature
@@
$\begin{split}\mathbb{E}\left[\mathcal{I}_{k}\right]=&\int\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{k},\bm{\Sigma}_{k})f(\tilde{\mathbf{x}}\mid\textbf{u})\psi(\textbf{u})\mathrm{d}\tilde{\mathbf{x}}\mathrm{d}\textbf{u}\\ =&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{k},\bm{\Sigma}_{k})\mu_{\bm{\psi}}(\tilde{\mathbf{x}})\mathrm{d}\tilde{\mathbf{x}},\\ \end{split}$ (22)
@@ 3.2 Inducing points and hyperparameter selection
Inducing point location
@@
For known GP hyperparameters, Burt et al. (2020) show that sampling from an $M$-determinantal point process ($M$-DPP) to select inducing points $\mathbf{Z}\subset\mathbf{X}$ will make the trace error term of the GP-elbo (Eq. 24) close to its optimal value. Since the DPP approach is intractable, we follow Burt et al. (2020) and Maddox et al. (2021) who recommend instead to sequentially select inducing points that greedily maximize the diagonal of the error term,
@@
GP hyperparameters
The procedure for the selection of inducing points mentioned above requires known GP hyperparameters $\bm{\xi}$. In practice, it is often enough to start with a reasonable estimate for $\bm{\xi}$ and then iterate the process multiple times (Burt et al., 2020). To obtain an initial estimate for the GP hyperparameters, we fit an exact GP on a small subset of the data and use the exact GP hyperparameters for initial inducing point selection. The subset is chosen via stratified $K$-means, where we ensure that the chosen subset is representative of the full set in terms of both location and log-density values. After the above initialization, we iterate sparse GP hyperparameter optimization and inducing points selection using the current sparse GP hyperparameters, until no improvement on the GP-elbo is found. Similarly to other block-optimization procedures, this process is not guaranteed to find the global optimum of the GP-elbo, but it often works well in practice (Burt et al., 2020).